OpenStudy (anonymous):

It is the year Googleplex and the whole world is voting for 1 President. Electoral College votes are distributed by the fraction of the world’s population. If China has 1/3 of the world’s population and India has 3/8 of the world’s population, and there are 2400 Electoral College votes, then how many of these votes are given to China and India?

OpenStudy (anonymous):

Cheena: 1/3, Rest of world is 2/3. India is 3/8 * 2/3 =6/24 [(1/3)*X]/100 + [(6/24)*X]/100 = Y Y * 100 is the percentage. Then you take the percentage multiply by 2400, to get your answer.